Fifteen years from now Mr. Thomas will be twice as old as Mrs. Thomas is now; whereas Mrs. Thomas will then be as old as Mr. Thomas will be 10 years from now. How old are Mr. and Mrs. Thomas now
Answer by stanbon(75887)

Fifteen years from now Mr. Thomas will be twice as old as Mrs. Thomas is now; whereas Mrs. Thomas will then be as old as Mr. Thomas will be 10 years from now. How old are Mr. and Mrs. Thomas now
----
Let Mr. be R
Let Mrs. be S
--------
(R+15) = 2(S)
(S+15) = R+10
----------------------
R = 2S - 15
Substitute for R and solve for "S":
S+15 = 2S-15+10
S+15 = 2S-5
S = 20 yrs (Mrs. Thomas's age now)
Since R = 2S-15, R = 2*20-15, R = 25 yrs (Mr. Thomas's age now)
--------------------------------------------------
Cheers,
Stan H.